OpenStudy (anonymous):

Solve by substitution method: 3x+9y=-6 -8+y=66

OpenStudy (anonymous):

-8+y = 66 Therefore y = 74 3x +9y = -6 to x +3y = -6 x = -3y-6 = -228

OpenStudy (anonymous):

Thank you